2(2x + 1) = x + 15 + 2x
Answer by sdmmadam@yahoo.com(530)

2(2x + 1) = x + 15 + 2x
4x+2 = 15 +(x+2x)
[expanding on the LHs, that is multiplying every term in the bracket by the 2 outside and on the RHS,using additive commutativity and associativity namely
x + 15 + 2x= (x+15)+2x = (15+x)+2x = 15 + (x+2x) which amounts to grouping like terms]
4x+2 = 15+3x
4x-3x=15-2 (grouping like terms by transferring terms from one side to another side changing sign whenever changing side. ) (or here in other words adding
(-3x) to both the sides and adding (-2) to both the sides )
x=13